OpenStudy (anonymous):

A circle with area A = pi r^2 is expanding with time. If dA/dt = 5cm^2/s, what is dr/dt when r=6?

OpenStudy (anonymous):

Hint:\[\frac{dA}{dt} = \frac{d}{dt}A = \frac{d}{dt}(\pi r^2)\]

OpenStudy (anonymous):

so it would be 2r?

OpenStudy (anonymous):

no, 2pi r dr/dt. (chain rule)

OpenStudy (anonymous):

so now what?

OpenStudy (anonymous):

\[\frac{dA}{dt} = \pi \frac{d}{dt}r^2 = \pi \frac{dr}{dt}(\frac{d}{dr}r^2) = \pi \frac{dr}{dt}2r\]Do you understand up to this point?

OpenStudy (ranga):

\[A = \pi r^2\\\frac{ dA }{ dt } = \frac{ d }{ dt }(\pi r^2) = 2\pi r \frac{ dr }{ dt }\]

OpenStudy (ranga):

Just plug in the numbers: dA/dt = 5, r = 6, find dr/dt

OpenStudy (anonymous):

how would I find dr and dt?

OpenStudy (anonymous):

Do you know what differentiation is?

OpenStudy (anonymous):

sort of

OpenStudy (anonymous):

It seems to me like maybe you get how to actually differentiate but you don't understand what exactly the result you're getting means. Something like dx/dt describes the relationship between two variables; it's not just an actual fraction that you can/would want to split up (in this context at least). For this particular example, dr/dt is the rate of change of the radius with respect to time, so all you have to do is evaluate "dr/dt" when r=6 if dA/dt = 5.